# Pulsed Vector Atomic Magnetometer Using an Alternating Fast-Rotating Field

### v3

@@ Abstract
We present a vector atomic magnetometer based on applying a fast-rotating magnetic field to a pulsed 87Rb scalar atomic magnetometer. This method enables simultaneous measurements of the total magnetic field and two polar angles relative to the plane of magnetic field rotation. Using two channels in a gradiometer mode, it provides simultaneous measurements of the total field gradient with a sensitivity of 35 $\mathrm{fT/\sqrt{Hz}}$ (0.7 part per billion), as well as two polar angles with resolutions of 6 $\mathrm{nrad/\sqrt{Hz}}$ at 50 $\mu$T Earth field strength. The noise spectrums of these measurements are flat down to 1 Hz and 0.1 Hz, respectively. Crucially, this approach avoids several metrological difficulties associated with vector magnetometers and gradiometers. We detail the fundamental, systematic, and practical limits of such vector magnetometers. Notably, we provide a comprehensive study of the systematic effects of vector atomic magnetometers. We introduce a new concept of dynamic heading error and investigate several other systematic effects. A unique peak-altering fast rotating field modulation is proposed to cancel out these systematics. Additionally, we derive fundamental limits on the sensitivity of such sensors and demonstrate that the vector sensitivity of the sensor can approach its scalar sensitivity while retaining the accuracy and metrological advantages of scalar sensors. This high-dynamic-range vector magnetometer, with ultrahigh resolution and inherent calibration, is suitable for a wide array of applications.
high dynamic range, vector atomic magnetometry, dynamic heading error, systematics
preprint: APS/123-QEP
I Introduction
@@
II Results
II.1 Spin dynamics
@@
As shown in Fig. 1a, we assume the spins are fully polarized along the negative y-axis. The magnetic field initially points along the z-axis, while a rotating field originates from the x-axis and rotates counterclockwise in the xy-plane. We consider the behavior of polarization in a magnetic field, where it undergoes precession in response to the applied field. This magnetic field is composed of both a static and a rotating component. In the presence of these components, the polarization vector evolves according to specific rotations. The spin evolution, without considering spin relaxations, can be expressed as:
$\mathbf{P}(t)=\mathcal{R}[\theta,\hat{z}]\cdot\mathcal{R}[\psi,\mathbf{B_{tot}}]\cdot(-\hat{y}),$ (2)
where $\mathcal{R}[\phi,\mathbf{v}]$ is a 3D rotation matrix for an anti-clockwise rotation of $\phi$ degrees around the vector $\mathbf{v}$. For the applied rotating field in the xy-plane, we have $B_{x}=B_{m}\sin(\omega_{m}t+\phi_{x})$, $B_{y}=B_{m}\sin(\omega_{m}t+\phi_{y})$, and $|\phi_{x}-\phi_{y}|=\pi/2$. Therefore, $\mathbf{B_{tot}}=(B_{m},0,B_{z}-\omega_{m}/\gamma)$. Here, $\hat{x}$, $\hat{y}$ and $\hat{z}$ are the unit vectors along the x, y and z axes, respectively. We define $\theta=\omega_{m}t$ and $\psi=\gamma t|\mathbf{B_{tot}}|$. Eventually, the spin projections can be written as:
@@
where $\omega_{0}=\gamma|\mathbf{B_{tot}}|$. By inserting residual transverse magnetic fields $b_{x}$ and $b_{y}$ into Eq. 2, we can get
@@
where $|\mathbf{B_{tot}}|=\sqrt{B_{m}^{2}+b_{x}^{2}+b_{y}^{2}+(B_{z}-\omega_{m}/\gamma)^{2}}$. The residual magnetic fields in the x-direction, $b_{x}$, and y-direction, $b_{y}$, are proportional to the amplitudes of the first harmonic, $\sin\omega_{m}t$ and $\cos\omega_{m}t$, respectively, in the phase shift. The conversion factor for converting from the amplitude of the phase shift to the transverse magnetic field is given by ±$\pm\omega_{m}\sqrt{B_{m}^{2}+B_{z}^{2}}/{\gamma B_{m}}$ (Table 1 details the sign dependence). Thus, by analyzing the phase shifts of the Larmor precession signal under the fast rotating field, we can measure both the transverse magnetic fields and the total magnetic field.
The intuitive classical model underlying the operating principle of the FRF vector magnetometer is illustrated in Fig. 1a. The scalar magnitude of the total magnetic field, $|\mathbf{B_{tot}}|$, depends on the angle between the rotating field $\mathbf{B_{m}}$ and the residual magnetic field $\mathbf{B_{0}}$. When a small residual transverse magnetic field $\mathbf{\Delta B}$—comprising $\mathbf{b_{x}}$ and $\mathbf{b_{y}}$—is present, $|\mathbf{B_{tot}}|$ varies with oscillating components at the rotation frequency of $\mathbf{B_{m}}$. The phases and amplitudes of these oscillations are influenced by the angle between $\mathbf{B_{0}}$ and $\mathbf{B_{m}}$. Consequently, the Larmor precession frequency, proportional to $|\mathbf{B_{tot}}|$, also displays oscillating components: a residual field in the x-direction, $b_{x}$, produces an in-phase oscillation, while a residual field in the y-direction induces an out-of-phase component.
@@
a) Experimental setup
c) Phase shift analysis
b) Four-shot rotating field scheme
Figure 1: (a) Three coils determine the x, y, and z directions. The sensor head sits on a rotating stage with the cell positioned at the stage’s center. The sensor head can freely rotate in the xz and yz-planes. The angle between the sensor head’s central axis and the z-axis is denoted as $\alpha$. $\beta$ is the angle by which the pump beam moves away from the position where the pump laser is perpendicular to the magnetic field. $\theta$ represents angle between the axis of rotation and the rotating field vector. Notably, WP stands for Wollaston prism, and Q-PD stands for quadrant photodiode. (b) Trig signal, when high after a rising edge, indicates the preparation phase: the AC heater heats the cell, and the pump laser polarizes the atoms. When low after a falling edge, the pump laser and AC heater are turned off, and the probe beam laser measures the FID signal. There are four shots with different start phases of $B_{x}$ and $B_{y}$.The start phases of the rotating field play a significant role; therefore, we indicate them with vertical dashed lines. This four-shot scheme is specifically developed to cancel out the systematic effects, such as Berry’s phase shift, dynamic heading error, probe beam heading error, eddy current, and the systematic caused by the threshold voltage of the MDA. The magnetic field only alters during the peak values (“Peak-Altering”) to suppress the eddy current caused by the rotating field. The signals during the preparation time are blanked out. (c) Upper: When a magnetic field of -87 nT is applied in the y-direction, the phase shifts of the FID signal acquired in b) are analyzed by the MDA as a function of time. Lower: when the magnetic fields in the x and y directions are roughly compensated. The first harmonic signals are suppressed and the second harmonic signal can be observed.
II.2 Systematics
@@
[Berry’s phase shift] The Berry phase shift refers to the geometric phase acquired by a spin system during an adiabatic, cyclic evolution [35]. As the magnetic field rotates slowly, the spins adiabatically follow the field. At the end of a complete cycle, when the field returns to its original configuration, the spins acquire a phase shift. If the system undergoes cyclic adiabatic evolution at a constant frequency, the Berry phase accumulates linearly over time. Thus, the Berry phase term as a function of time, from Eq. 3, can be written as
@@
where $\theta\approx B_{m}/B_{z}$ is the angle between the total magnetic field and z-axis as shown in Fig. 1a. The Berry’s phase acquired over one complete period, given by Eq. 6 as $2\pi(1-\cos{\theta})$, is consistent with Berry’s phase shift of Nuclear Magnetic Resonance (NMR) described in Ref. [35].
[Second harmonic phase shift] Another phase shift from Eq. 3 is proportional to $\sin(2\omega_{m}t)$, which can be written as
@@
[Heading errors] The heading error of atomic magnetometers refers to the dependence of the measured magnetic field values on the orientation of the sensor relative to the magnetic field [36]. The pump beam heading error is well-studied and is primarily caused by nonlinear Zeeman splitting and the difference between Zeeman resonances in the two hyperfine ground states. We set nuclear Landé factor $g_{I}\approx 0$ [37] and the heading error can be well described by $B_{SH}=\mathcal{B_{H}}\sin\beta$, where
@@
We discovered that the dynamic heading error can be explained by the fact that the spin precession plane adiabatically follows the rotation of the total magnetic field. This maintains a constant relative angle between the total magnetic field vector and the spin precession plane, resulting in a stable precession frequency, as the weighted sum of different Zeeman transitions remains consistent. Consequently, the dynamic heading error depends on the initial angle between the spin polarization and the total magnetic field, which remains constant and equal to its initial value. (More details in Sec.  II.3).
To clarify the elimination of the systematic dynamic heading error, we separate the dynamic heading error, $B_{DH}$, into two components: a static part ($B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta$), which is equivalent to the heading error in the absence of the rotating field ($\approx\mathcal{B_{H}}\sin\beta$) when $\theta$ is small, and a rotating-field-phase-dependent part ($B_{PD}$). Notably, the static heading error is independent of the rotating field phases.
There’s an analogous “heading error” effect from the probe beam. This error creates a systematic effect where the measured transverse magnetic field depends on the relative angle between the probe beam and the plane of the rotating field, analogous to the pump beam heading error. Regarding the probe beam heading error, we define the angle that the probe beam rotates away from the x-axis as $\alpha$, with the sensor rotating in the xz-plane., as shown in Fig. 1(b). The spin projection along the probe beam can be written as $P_{prob}(t)=P_{x}(t)\cos\alpha+P_{z}(t)\sin\alpha$. By inserting Eqs. 3 and 4, and analyzing the first harmonic phase shift that depends on $\alpha$, we can determine the phase shift caused by the probe beam heading error
@@
[Eddy current] Another systematic error is induced by eddy currents. The altering rotating field can generate eddy currents on the electrically conductive aluminum or $\mu$-metal magnetic shield, which in turn generates a magnetic field. The magnitude of the eddy current magnetic field highly depends on how the rotating magnetic field is altered. (More details in Sec. II.3)
[Sign dependence of systematics] We discovered that the signs of these systematic errors depend on the phases of the rotating field, while their absolute magnitudes remain unchanged. This characteristic offers an opportunity to eliminate these systematics. The sign of the $B_{PD}$ part of dynamic heading error depends on the rotating field’s phases, but the sign of the static heading error does not. We introduce a four-shot scheme (shown in Fig. 1b) by altering the phases of the rotating field in the order listed in Table 1. By averaging these four shots, the systematic errors can be canceled out.
@@ II.3 Experimental results
As shown in Fig. 1a, a rotating magnetic field with an amplitude of approximately 18 $\mathrm{\mu T}$ and a freuqency of $\omega_{m}=2\pi\times 480$ Hz is applied in the transverse plane, and a leading magnetic field $B_{z}$ is applied along the longitudinal direction. The total amplitude of the applied magnetic field is maintained at approximately 50 $\mathrm{\mu T}$. The frequency of the pump beam pulse is set to 348 kHz, synchronized with the Larmor frequency to achieve the highest pumping rate. The duty cycle of a pump beam pulse is approximately 1.4%.
@@
The phase shift of the acquired signal relative to the Larmor frequency is analyzed using the HP 53310A modulation domain analyzer (MDA), which detects the zero-crossings of the precession signal and a reference signal of $\omega_{\text{ref}}$ that is close to the precession frequency. The phase shifts between the zero crossings of the precession signal and the reference signal are then calculated. The phase shift measuremment of one block (four shots) of data is shown in Fig. 1c.
The fitting model further fits the measured phase shifts $\delta\Phi(t)$ from the MDA as a function of measurement time to get: offset, slope ($\Phi_{s}$, Table A1 in the supplemental material details the components of the slope), amplitudes of the first harmonics ($b_{x}\propto\Phi_{x}$ and $b_{y}\propto\Phi_{y}$), second harmonics of $\omega_{m}$, and amplitudes of the first harmonics of the hyperfine frequency $\omega_{\text{hp}}$ (the difference in precession frequency for the F=1 and F=2 atoms induces an exponentially decaying sinusoidal phase shift at a frequency of $\omega_{\text{hp}}=2\pi\times 1.39$ kHz, see supplemental material Sec. I).
@@
II.3.1 Results of the systematics
The experimental results for the systematics are presented in Table 2.

### v4

@@ Abstract
We introduce a vector atomic magnetometer that employs a fast-rotating magnetic field applied to a pulsed 87Rb scalar atomic magnetometer. This approach enables simultaneous measurements of the total magnetic field and its two polar angles relative to the rotation plane. Operating in gradiometer mode, the magnetometer achieves a total field gradient sensitivity of 35 $\mathrm{fT/\sqrt{Hz}}$ (0.7 parts per billion) and angular resolutions of 6 $\mathrm{nrad/\sqrt{Hz}}$ at a 50 $\mu$T Earth field strength. The noise spectra remain flat down to 1 Hz and 0.1 Hz, respectively. Here we show that this method overcomes several metrological challenges commonly faced by vector magnetometers and gradiometers. We propose a unique peak-altering modulation technique to mitigate systematic effects, including a newly identified dynamic heading error. Additionally, we establish the fundamental sensitivity limits of the sensor, demonstrating that its vector sensitivity approaches scalar sensitivity while preserving the inherent accuracy and calibration benefits of scalar sensors. This high-dynamic-range, ultrahigh-resolution magnetometer offers exceptional versatility for diverse applications.
1 Introduction
@@
2 Results
2.1 Spin dynamics
@@
As shown in Fig. 1, we assume the spins are fully polarized along the negative y-axis. The magnetic field initially points along the z-axis, while a rotating field originates from the x-axis and rotates counterclockwise in the xy-plane. We have $B_{x}=B_{m}\sin(\omega_{m}t+\phi_{x})$, $B_{y}=B_{m}\sin(\omega_{m}t+\phi_{y})$, and $|\phi_{x}-\phi_{y}|=\pi/2$. For this assumption, $\phi_{x}=\pi/2$ and $\phi_{y}=0$. We consider the behavior of polarization in a magnetic field, where it undergoes precession in response to the applied field. This magnetic field is composed of both a static and a rotating component. In the presence of these components, the polarization vector evolves according to specific rotations. The spin evolution, without considering spin relaxations, can be expressed as:
$\mathbf{P}(t)=\mathcal{R}[\theta,\hat{z}]\cdot\mathcal{R}[\psi,\mathbf{\tilde{B}_{tot}}]\cdot(-\hat{y}),$ (2)
where $\mathcal{R}[\phi,\mathbf{v}]$ is a 3D rotation matrix for an anti-clockwise rotation of $\phi$ degrees around the vector $\mathbf{v}$. In the co-rotating reference frame, the effective field after applying the rotating wave approximation is given by $\mathbf{\tilde{B}_{tot}}=(B_{m},0,B_{z}-\omega_{m}/\gamma)$. Here, $\hat{x}$, $\hat{y}$ and $\hat{z}$ are the unit vectors along the x, y and z axes, respectively. We define $\theta=\omega_{m}t$ and $\psi=\gamma t|\mathbf{\tilde{B}_{tot}}|$. Eventually, the spin projections can be written as:
@@
where $\omega_{0}=\gamma|\mathbf{\tilde{B}_{tot}}|$. By inserting residual transverse magnetic fields $b_{x}$ and $b_{y}$ into Eq. 2, we can get
@@
where $|\mathbf{\tilde{B}_{tot}}|=\sqrt{B_{m}^{2}+b_{x}^{2}+b_{y}^{2}+(B_{z}-\omega_{m}/\gamma)^{2}}$. The residual magnetic fields in the x-direction, $b_{x}$, and y-direction, $b_{y}$, are proportional to the amplitudes of the first harmonic, $\sin\omega_{m}t$ and $\cos\omega_{m}t$, respectively, in the phase shift. The conversion factor for converting from the amplitude of the phase shift to the transverse magnetic field is given by ±$\pm\omega_{m}\sqrt{B_{m}^{2}+B_{z}^{2}}/{\gamma B_{m}}$ (Table 1 details the sign dependence for $b_{x}$ and $b_{y}$ across various starting phases). Thus, by analyzing the phase shifts of the Larmor precession signal under the fast rotating field, we can measure both the transverse magnetic fields and the total magnetic field.
The intuitive classical model underlying the operating principle of the FRF vector magnetometer is illustrated in Fig. 1. The scalar magnitude of the total magnetic field, $|\mathbf{B_{tot}}|$, depends on the angle between the rotating field $\mathbf{B_{m}}$ and the residual magnetic field $\mathbf{B_{0}}$. When a small residual transverse magnetic field $\mathbf{\Delta b}$—comprising $\mathbf{b_{x}}$ and $\mathbf{b_{y}}$—is present, $|\mathbf{B_{tot}}|$ varies with oscillating components at the rotation frequency of $\mathbf{B_{m}}$. The phases and amplitudes of these oscillations are influenced by the angle between $\mathbf{B_{0}}$ and $\mathbf{B_{m}}$. Consequently, the Larmor precession frequency, proportional to $|\mathbf{B_{tot}}|$, also displays oscillating components: a residual field in the x-direction, $b_{x}$, produces an ou-of-phase oscillation, while a residual field in the y-direction induces an in-phase component.
Figure 1: Experimental setup: Three coils determine the x, y, and z directions. The sensor head sits on a rotating stage with the cell positioned at the stage’s center. The sensor head can freely rotate in the xz and yz-planes. $\alpha$ is the angle by which the probe beam moves away from the position where the probe beam is perpendicular to the static magnetic field in xz-plane. $\beta$ is the angle by which the pump beam moves away from the position where the pump beam is perpendicular to the static magnetic field in yz-plane. $\theta$ represents angle between the axis of rotation and the rotating field vector. Notably, WP stands for Wollaston prism, and Q-PD stands for quadrant photodiode.
Figure 2: Four-shot rotating field scheme: Trig signal, when high after a rising edge, indicates the preparation phase: the AC heater heats the cell, and the pump laser polarizes the atoms. When low after a falling edge, the pump laser and AC heater are turned off, and the probe beam laser measures the FID signal. There are four shots with different start phases of $B_{x}$ and $B_{y}$. The start phases of the rotating field play a significant role as shown in Table 1; therefore, we indicate them with vertical dashed lines. This four-shot scheme is specifically developed to cancel out the systematic effects, such as Berry’s phase shift, dynamic heading error, probe beam heading error, eddy current, and the systematic caused by the threshold voltage of the MDA. The magnetic field only alters during the peak values (“Peak-Altering”) to suppress the eddy current caused by the rotating field. The signals during the preparation time are blanked out.
Figure 3: Phase shift analysis: Upper: when a magnetic field of -87 nT is applied in the y-direction, the phase shifts of the FID signal acquired in Fig. 2 are analyzed by the MDA as a function of time. Lower: when the magnetic fields in the x and y directions are roughly compensated. The first harmonic signals are suppressed and the second harmonic signal can be observed.
@@
2.2 Systematics
@@
2.2.1 Berry’s phase shift
The Berry phase shift refers to the geometric phase acquired by a spin system during an adiabatic, cyclic evolution [35]. As the magnetic field rotates slowly, the spins adiabatically follow the field. At the end of a complete cycle, when the field returns to its original configuration, the spins acquire a phase shift. If the system undergoes cyclic adiabatic evolution at a constant frequency, the Berry phase accumulates linearly over time. Thus, the Berry phase term as a function of time, from Eq. 3, can be written as
@@ 2.2.1 Berry’s phase shift
where $\theta\approx B_{m}/B_{z}$ is the angle between the total magnetic field and z-axis as shown in Fig. 1. The Berry’s phase acquired over one complete period, given by Eq. 6 as $2\pi(1-\cos{\theta})$, is consistent with Berry’s phase shift of Nuclear Magnetic Resonance (NMR) described in Ref. [35].
2.2.2 Second harmonic phase shift
Another phase shift from Eq. 3 is proportional to $\sin(2\omega_{m}t)$, which can be written as
@@ 2.2.2 Second harmonic phase shift
2.2.3 Heading errors
The heading error of atomic magnetometers refers to the dependence of the measured magnetic field values on the orientation of the sensor relative to the magnetic field [36]. The pump beam heading error is well-studied and is primarily caused by nonlinear Zeeman splitting and the difference between Zeeman resonances in the two hyperfine ground states. We set nuclear Landé factor $g_{I}\approx 0$ [37] and the heading error can be well described by $B_{S\!H}=\mathcal{B_{H}}\sin\beta$, where
@@ 2.2.3 Heading errors
We discovered that the dynamic heading error can be explained by the fact that the spin precession plane adiabatically follows the rotation of the total magnetic field. This maintains a constant relative angle between the total magnetic field vector and the spin precession plane, resulting in a stable precession frequency, as the weighted sum of different Zeeman transitions remains consistent. Consequently, the dynamic heading error depends on the initial angle between the spin polarization and the total magnetic field, which remains constant and equal to its initial value. (More details in Sec. 2.3).
To clarify the elimination of the systematic dynamic heading error, we separate the conventional dynamic heading error, $B_{D\!H}$, into two components: a static part ($B_{S\!H}=\mathcal{B_{H}}\cos\theta\sin\beta$), which is equivalent to the heading error in the absence of the rotating field ($\approx\mathcal{B_{H}}\sin\beta$) when $\theta$ is small, and a rotating-field-phase-dependent part ($B_{P\!D}$). Notably, the static heading error is independent of the rotating field phases.
There’s an analogous “heading error” effect from the probe beam. This error creates a systematic effect where the measured transverse magnetic field depends on the relative angle between the probe beam and the plane of the rotating field, analogous to the pump beam heading error. Regarding the probe beam heading error, we define the angle that the probe beam rotates away from the x-axis as $\alpha$, with the sensor rotating in the xz-plane., as shown in Fig. 1. The spin projection along the probe beam can be written as $P_{prob}(t)=P_{x}(t)\cos\alpha+P_{z}(t)\sin\alpha$. By inserting Eqs. 3 and 4, and analyzing the first harmonic phase shift that depends on $\alpha$, we can determine the phase shift caused by the probe beam heading error
@@
2.2.4 Eddy current
Another systematic error is induced by eddy currents. The altering rotating field can generate eddy currents on the electrically conductive aluminum or $\mu$-metal magnetic shield, which in turn generates a magnetic field. The magnitude of the eddy current magnetic field highly depends on how the rotating magnetic field is altered. (More details in Sec. 2.3)
2.2.5 Sign dependence of systematics
We discovered that the signs of these systematic errors depend on the phases of the rotating field, while their absolute magnitudes remain unchanged. This characteristic offers an opportunity to eliminate these systematics. The sign of the $B_{P\!D}$ part of dynamic heading error depends on the rotating field’s phases, but the sign of the static heading error does not. We introduce a four-shot scheme (shown in Fig. 2) by altering the phases of the rotating field in the order listed in Table 1. By averaging these four shots, the systematic errors can be canceled out.
@@ 2.3 Experimental results
As shown in Fig. 1, a rotating magnetic field with an amplitude of approximately 18 $\mathrm{\mu T}$ and a freuqency of $\omega_{m}=2\pi\times 480$ Hz is applied in the transverse plane, and a leading magnetic field $B_{z}$ is applied along the longitudinal direction. The total amplitude of the applied magnetic field is maintained at approximately 50 $\mathrm{\mu T}$. The frequency of the pump beam pulse is set to 348 kHz, synchronized with the Larmor frequency to achieve the highest pumping rate. The duty cycle of a pump beam pulse is approximately 1.4%.
@@
The phase shift of the acquired signal relative to the Larmor frequency is analyzed using the HP 53310A modulation domain analyzer (MDA), which detects the zero-crossings of the precession signal and a reference signal of $\omega_{\text{ref}}$ that is close to the precession frequency. The phase shifts between the zero crossings of the precession signal and the reference signal are then calculated. The phase shift measuremment of one block (four shots) of data is shown in Fig. 3.
The fitting model further refines the measured phase shifts, $\delta\Phi(t)$, for each shot from the MDA, modeling them as a function of measurement time to obtain: offset, slope ($\Phi_{s}$, Supplementary Table 2 in the Supplemental Information details the components of the slope), amplitudes of the first harmonics ($b_{x}\propto\Phi_{x}$ and $b_{y}\propto\Phi_{y}$), second harmonics of $\omega_{m}$, and amplitudes of the first harmonics of the hyperfine frequency $\omega_{hp}$ (the difference in precession frequency for the F=1 and F=2 atoms induces an exponentially decaying sinusoidal phase shift at a frequency of $\omega_{hp}=2\pi\times 1.39$ kHz, see Supplementary Note 1 in the Supplemental Information).
@@
2.3.1 Results of the systematics
The experimental results for the systematics are presented in Table 2. The Berry phase will result in a phase shift that adds to the total magnetic field measurement. When $\theta\approx 21^{\circ}$, magnetic field measurement systematic caused by Berry’s phase shift is measured to be 5.1 nT, while Eq. 6 predicts 4.7 nT ($\phi_{B}/\gamma$), as shown in Table. 2.